Calculate the wavelength in meters of electromagnetic radiation that has a frequency of 980.0 kHz. (c = 3.00 X 10⁸ m/s)


1
Expert's answer
2021-03-08T05:56:15-0500

V= 980.0 x 103Hz

C= 3.00 x 108m/s

¥= ?

C= V¥

¥= C/V = (3.00 x 108)/(980.0 x 103= 306.12m